The length of human pregnancies from conception to birth varies according to a distribution that is approximately Normal with mean 272 days and standard deviation 16 days. 95% of all pregnancies last between
Answer by edjones(8007)

1.96 and -1.96 are the 95% bounds.
(x-272)/16=-1.96
x-272=-31.36
x=240.6
.
(x-272)/16=1.96
x-272=31.96
x=304
.
Between 240.6 and 304 days
.
Ed